# Non-Equilibrium Steady States and Quantum Chaos in a three-site Driven-Dissipative Bose-Hubbard Chains base on Self-Consistent Mean-Field Approach

### v5

@@ V OTOC
Figure 2: (Left) Time evolution of the local Phase OTOC $D_{3,3}(t)$ calculated on NESS background using the Picard iteration self-consistent solver. The rapid initial growth and subsequent saturation at a non-zero value indicate significant information scrambling and the onset of local quantum chaos induced by the Kerr nonlinearity. (Right) The OTOC solved by exact solver.
@@ VI Effective Liouvillian $\mathbf{L}_{eff}$
where the cross terms $\delta\hat{a}(t),\hat{\mathcal{F}}^{\dagger}(t)]$ vanish because initial operators commute with future noise operators). According to fluctuation-dissipation theorem, this implies $\hat{\mathcal{F}}(t),\hat{\mathcal{F}}^{\dagger}(t)]=1-e^{-\gamma t}$. The accumulated noise $\hat{\mathcal{F}}(t)$ exactly replenishes the commutator as it decays. Without $\hat{\mathcal{F}}(t)$, the commutator would go to zero, which is a violation of the laws of quantum mechanics.
@@
Since $\det(e^{\mathbf{A}})=e^{\text{Tr}(\mathbf{A})}$ and $\text{Tr}(\mathbf{L}_{\text{eff}})=(-i\tilde{\Delta}-\frac{\gamma}{2})+(i\tilde{\Delta}-\frac{\gamma}{2})=-\gamma$, we have $\det(e^{\mathbf{L}_{\text{eff}}t})=u(t)u^{*}(t)-v(t)v^{*}(t)=|u(t)|^{2}-|v(t)|^{2}=e^{\text{Tr}(\mathbf{L}_{\text{eff}})t}=e^{-\gamma t}$. Using Cayley-Hamilton theorem, the 2-by-2 matrix $\mathbf{M}:=\mathbf{L}_{eff}+\frac{\gamma}{2}{\bf I}$ satisfy $\mathbf{M}^{2}-\text{tr}(\mathbf{M})\mathbf{M}+\det(\mathbf{M})\mathbf{I}=\mathbf{0}$, thus $\mathbf{M}^{2}=-\det(\mathbf{M})\mathbf{I}=-\Omega^{2}{\bf I}$. Note that in oscillatory stable regime, $\det(\mathbf{M})=(-i(\Delta+2U|\Psi_{SS}|^{2}))(i(\Delta+2U|\Psi_{SS}|^{2}))-(-iU\Psi_{SS}^{2})(iU(\Psi_{SS}^{*})^{2})=(\Delta+2U|\Psi_{SS}|^{2})^{2}-|U\Psi_{SS}^{2}|^{2}=\Omega^{2}\geq 0$.
@@ VII Truncated Wigner Approximation (TWA)
For a symmetric vacuum state, the variance of the complex variable is $\langle|\alpha|^{2}\rangle=\langle X^{2}\rangle+\langle P^{2}\rangle=1/2$, with $\langle X^{2}\rangle=\langle P^{2}\rangle=1/4$ and $\Delta X=\Delta P=1/2$ (which is possibel only for the vacuum or a coherent/unsqueezed state). The product of uncertainties in linear system with Gaussian state is $\Delta X\Delta P=\frac{1}{4}$, corresponds to the minimal uncertainty state and minimal variance ($1/2$) in TWA. This is consistent with above symmetric ordering requirement $\overline{|\alpha|^{2}}=\langle\hat{n}\rangle+1/2$. The noise $\sqrt{\gamma/2}dW$ which inject randomness is isotropic while the nonlinearity ($U$) is anisotropic and cause non-Gaussian state with $\Delta X\Delta P>\frac{1}{4}$. As the system relaxes from the vacuum to a driven NESS, the total variance $|\alpha|^{2}$ increases due to the fluctuation part of TWA which suppressing the coherent amplitude and restoring the $U(1)$ symmetry. While the sum of variances is bounded by the noise term, the individual quadrature variance (e.g., $\Delta X^{2}$) can be compressed below 1/4.
The Kerr term $\hat{n}(\hat{n}-1)$ maps to the Wigner symbol $H_{int}=\frac{U}{2}(|\alpha|^{4}-2|\alpha|^{2}+1/2)$. The drift $\mathcal{A}$ in Langevin equation is derived from the derivative (Poisson bracket) $i\frac{\partial H_{int}}{\partial\alpha^{*}}$,
@@
where the quantum bias 1-1 is related to the above bias 1/2 in the operator $\alpha$ in symmetric (Weyl) ordering and ensures the classical trajectories obeys the quantum commutation relations during evolution.
@@
Note that in any cases, we consider the deturning term $\propto\Delta n<0$ (i.e., $i\Delta\alpha$ in Heisenberg equation of motion), in consistent with the content in above sections. From the Heisenberg equation $\dot{\hat{a}}=-i[\hat{a},\hat{H}$], the interaction part is $i[\hat{a},\frac{U}{2}\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}]=-iU\hat{a}^{\dagger}\hat{a}\hat{a}=-iU\hat{n}\hat{a}$. The TWA variables $\alpha$ in symmetric (Weyl) ordering can be used to map the operator product $\hat{n}\hat{a}=\hat{a}^{\dagger}\hat{a}\hat{a}$ to phase space, by the Weyl symbol of $\hat{a}^{\dagger}\hat{a}\hat{a}$ $|\alpha|^{2}\alpha-\alpha=(|\alpha|^{2}-1)\alpha$, where 1-1 is a quantum correction (ordering bias) that accounts for the commutation relation $\hat{a},\hat{a}^{\dagger}]=1$. The term $|\alpha|^{2}\alpha$ represents the classical mean-field interaction. The term $\alpha$ represents the self-interaction with vacuum noise that subtracts the over-counted zero-point energy inherent in the Wigner representation. The hopping term $J(\alpha_{l-1}+\alpha_{l+1})$ describes the coherent exchange of particles between neighboring sites in the chain. The dissipative damping $\frac{\gamma_{l}}{2}\alpha_{l}$ comes from the $\mathcal{D}[\sqrt{\gamma}\hat{a}$] term in the Lindblad equation. It causes the amplitude to decay exponentially toward zero in the absence of a drive. The stochastic term represents the vacuum noise entering through the loss channels, with the Markovian white noise $\langle d\mathcal{W}_{l}(t)d\mathcal{W}_{k}^{*}(t^{\prime})\rangle=\delta_{lk}\delta(t-t^{\prime})dt$. Since the loss is local to each site, $\mathcal{B}$ is a diagonal matrix with $\mathcal{B}_{ll}=\sqrt{\gamma_{l}/2}$ and scales the random noise by the square root of the dissipation rate. The complex wiener process $d\mathcal{W}_{l}$ represents independent random variables sampled at each time step. The evolution of the fluctuation spinor $\mathbf{v}=(\delta\hat{a},\delta\hat{a}^{\dagger})^{T}$ is governed by the linearized equation $\dot{\mathbf{v}}=\mathbf{L}_{\text{eff}}\mathbf{v}$. The matrix $\mathbf{L}_{\text{eff}}$ is defined as the Jacobian matrix of the total deterministic drift function $\dot{\alpha}=\mathcal{A}(\alpha,\alpha^{*})$,
@@
Figure 4: (a) The chaotic regime ($U=2.5,F=6,J=0.8$) where nonlinearity introduces phase scrambling. At site 1, the drive $F$ is strong enough to maintain a coherent offset. At sites 2 and 3, the scrambling $U$ is stronger than the transfer $J$. This generates diffusion and randomizes the phase, and the Wigner distribution is centered at (0,0), effectively restoring the $U(1)$ symmetry. (b) The linear regime ($U=0,F=1,J=0.8$) where all sites are displaced from the origin. The phase is pinned and U(1) symmetry is broken by the drive because the coherence $J$ is dominate over the randomizing force.

### v6

@@ V OTOC
A large OTOC implies that information has been scrambled before reaching the drain site because it measures the growth of operator complexity and the spreading of local perturbations across the system’s degrees of freedom. In our model, the OTOC specifically tracks the non-commutativity of operators at different sites, which quantifies how an initial local change at the first site evolves into a global change that affects the third site. While high dissipation at the third site reduces the total particle density by acting as a sink, it does not prevent the butterfly effect driven by the internal Hamiltonian’s nonlinearities and hopping amplitudes. If the internal scrambling rate is high, the information regarding the system’s initial state is hidden within complex, multi-particle correlations before the particles themselves are removed by the environment. Therefore, a large OTOC indicates that even though there are fewer particles reaching the drain, the phase or memory of the input signal has already been randomized or lost to local probes due to the chaotic nature of the dynamics, meaning the scrambling process is faster and more dominant than the dissipative loss. This ensures that the information is rendered inaccessible before it can be leaked out as coherent data, effectively signaling that the chaotic nature of the state is established regardless of the overall decay in particle number.
As shown in Fig.2, for both the GMF approximation and exact solver, the OTOC decays to zero at long time. In stable systems, any initial perturbation (the "kick" from $\hat{a}_{j}^{\dagger}$) is eventually damped by $\gamma$. The sensitivity $\partial a(t)/\partial a(0)$ vanishes as $t\to\infty$. In contrast, a chaotic system would show the OTOC saturating at a large value ($\approx 2n^{2}$) due to phase randomization.
An alternative expression of OTOC is related to the semiclassical Jacobian
$D_{1,L}(\tau)=\left|\frac{\partial\alpha_{L}(\tau)}{\partial\alpha_{1}(0)}\right|^{2}\approx|G^{R}_{L,1}(\tau)|^{2}=\left|\left(e^{\mathbf{L}_{\text{chain}}\tau}\right)_{L,1}\right|^{2},$ (48)
where the coupled chain Jacobian (a $2L\times 2L$ matrix) includes the effect of coupling $J$. In TWA, the deterministic evolution is given by $\dot{\alpha}_{l}=\mathcal{A}_{l}(\{\alpha\})$ with complex amplitudes $\alpha_{l}$. Let $\mathbf{V}(t)=(\delta\alpha_{1},\delta\alpha_{1}^{*},\dots,\delta\alpha_{L},\delta\alpha_{L}^{*})^{T}$ be the fluctuation vector. Linearizing around the steady state
$\frac{d\mathbf{V}(t)}{dt}=\mathbf{J}\mathbf{V},$ (49)
where $J_{ij}=\frac{\partial\dot{V}_{i}}{\partial V_{j}}$ and the solution is $\mathbf{V}(t)=e^{\mathbf{J}t}\mathbf{V}(0)$. The sensitivity of site $L$ to a perturbation at site 1 is
$\frac{\partial\alpha_{L}(\tau)}{\partial\alpha_{1}(0)}=\left(e^{\mathbf{J}\tau}\right)_{2L-1,1}$ (50)
with $\mathbf{J}$ the $2L\times 2L=6\times 6$ Jocobian matrix.
For the deterministic evolution of complex fields $\dot{\alpha}_{l}=\mathcal{A}_{l}(\alpha_{1},\alpha_{1}^{*},\dots,\alpha_{L},\alpha_{L}^{*})$, the steady state at equilibrium $\{\alpha_{SS}\}$ is defined as $\mathcal{A}_{l}(\{\alpha_{SS}\})=0\quad\text{for all }l$. We assume the system is slightly perturbed from the steady state $\alpha_{l}(t)=\alpha_{l,SS}+\delta\alpha_{l}(t)$, $\dot{\alpha}_{l}(t)=\frac{d}{dt}(\alpha_{l,SS}+\delta\alpha_{l}(t))=\delta\dot{\alpha}_{l}(t)$. Expand the function $\mathcal{A}_{l}$ around the steady state value using a first-order Taylor series $\mathcal{A}_{l}(\{\alpha\})\approx\mathcal{A}_{l}(\{\alpha_{SS}\})+\sum_{j}\left.\frac{\partial\mathcal{A}_{l}}{\partial\alpha_{j}}\right|_{SS}\delta\alpha_{j}+\sum_{j}\left.\frac{\partial\mathcal{A}_{l}}{\partial\alpha_{j}^{*}}\right|_{SS}\delta\alpha_{j}^{*}+\mathcal{O}(\delta^{2})$. Thus $\delta\dot{\alpha}_{l}\approx\sum_{j}\left(\frac{\partial\mathcal{A}_{l}}{\partial\alpha_{j}}\delta\alpha_{j}+\frac{\partial\mathcal{A}_{l}}{\partial\alpha_{j}^{*}}\delta\alpha_{j}^{*}\right)$. We define the column vector $\mathbf{V}$ as the collection of all fluctuations and their conjugates $\mathbf{V}=(\delta\alpha_{1},\delta\alpha_{1}^{*},\dots,\delta\alpha_{L},\delta\alpha_{L}^{*})^{T}$. The entries of the Jacobian matrix $\mathbf{J}$ are precisely the partial derivatives calculated at the steady state $J_{ij}=\frac{\partial\dot{V}_{i}}{\partial V_{j}}$. If $V_{i}=\delta\alpha_{l}$ and $V_{j}=\delta\alpha_{k}$, then $J_{ij}=\frac{\partial\mathcal{A}_{l}}{\partial\alpha_{k}}$. If $V_{i}=\delta\alpha_{l}$ and $V_{j}=\delta\alpha_{k}^{*}$, then $J_{ij}=\frac{\partial\mathcal{A}_{l}}{\partial\alpha_{k}^{*}}$.
In the TWA or GMF limit, the system is described by complex amplitudes $\alpha_{l}$. The sensitivity $D_{1,L}(\tau)=|\partial\alpha_{L}(\tau)/\partial\alpha_{1}(0)|^{2}$ is obtained from the $2L\times 2L$ Jacobian matrix $\mathbf{J}$,
$\mathbf{J}=\begin{pmatrix}\mathbf{M}_{1}&\mathbf{J}_{hop}&0\\ \mathbf{J}_{hop}&\mathbf{M}_{2}&\mathbf{J}_{hop}\\ 0&\mathbf{J}_{hop}&\mathbf{M}_{3}\end{pmatrix}$ (51)
where the local blocks $\mathbf{M}_{l}$ and hopping blocks $\mathbf{J}_{hop}$ are
$\mathbf{M}_{l}=\begin{pmatrix}-i(\Delta+2U|\Psi_{l}|^{2}-U)-\frac{\gamma_{l}}{2}&-iU\Psi_{l}^{2}\\ iU(\Psi_{l}^{*})^{2}&i(\Delta+2U|\Psi_{l}|^{2}-U)-\frac{\gamma_{l}}{2}\end{pmatrix},\quad\mathbf{J}_{hop}=\begin{pmatrix}iJ&0\\ 0&-iJ\end{pmatrix}$ (52)
The sensitivity is the element $(\exp(\mathbf{J}t))_{2L-1,1}$.
@@
Figure 2: Operator OTOC governed by the adjoint Liouvillian $e^{\mathcal{L}^{{\dagger}}t}$ where phase diffusion in a trajectory picture does not imply persistent operator growth in $\mathcal{L}^{{\dagger}}$ and dissipation dominate the long-time behavior. (Left) Time evolution of the local operator OTOC $D_{i,i}(t)$ calculated on NESS background using the Picard iteration self-consistent solver. Blue, green, and red correspond to site 1, site 2, and site 3, respectively. The real line and dashed line correspond to setting $F=2.5$ and $F=1$, respectively. Our parameters are setted as $J=1,\ U=2,\ \Delta=-0.5,\ \gamma_{1}=0.1=\gamma_{2}=0.1,\ \gamma_{3}=1$. The rapid initial growth and subsequent saturation at a non-zero value indicate significant information scrambling and the onset of local quantum chaos induced by the Kerr nonlinearity. (Right) Operator OTOC $D_{i,1}(t)$ solved by exact solver. Different to the above GMF approximation which assumes the total state is a product state $\hat{\rho}_{tot}=\hat{\rho}_{1}\otimes\hat{\rho}_{2}\otimes\hat{\rho}_{3}$ and thus forbids any entanglement or quantum correlations between sites and only captures local quantum effects (like onsite squeezing or photon blockade), the TWA allows trajectories of different sites to be correlated and captures classical-like correlations as well as the scrambling. However, it cannot represent purely non-local quantum entanglement (such as EPR pairs or Bell states) because it is still based on point-like trajectories in phase space. The exact solver acts on the full many-body Hilbert space and it captures every possible quantum correlation, including multi-site entanglement and non-local phase coherence.
@@ VI Effective Liouvillian $\mathbf{L}_{eff}$
where the cross terms $\delta\hat{a}(t),\hat{\mathcal{F}}^{\dagger}(t)]$ vanish because initial operators commute with future noise operators). According to fluctuation-dissipation theorem, this implies $\hat{\mathcal{F}}(t),\hat{\mathcal{F}}^{\dagger}(t)]=1-e^{-\gamma t}$. The accumulated noise $\hat{\mathcal{F}}(t)$ exactly replenishes the commutator as it decays. Without $\hat{\mathcal{F}}(t)$, the commutator would go to zero, which is a violation of the laws of quantum mechanics[18, 19].
@@
Since $\det(e^{\mathbf{A}})=e^{\text{Tr}(\mathbf{A})}$ and $\text{Tr}(\mathbf{L}_{\text{eff}})=(-i(\Delta+2U|\Psi_{SS}|^{2})-\frac{\gamma}{2})+(i(\Delta+2U|\Psi_{SS}|^{2})-\frac{\gamma}{2})=-\gamma$, we have $\det(e^{\mathbf{L}_{\text{eff}}t})=u(t)u^{*}(t)-v(t)v^{*}(t)=|u(t)|^{2}-|v(t)|^{2}=e^{\text{Tr}(\mathbf{L}_{\text{eff}})t}=e^{-\gamma t}$. Using Cayley-Hamilton theorem, the 2-by-2 matrix $\mathbf{M}:=\mathbf{L}_{eff}+\frac{\gamma}{2}{\bf I}$ satisfy $\mathbf{M}^{2}-\text{tr}(\mathbf{M})\mathbf{M}+\det(\mathbf{M})\mathbf{I}=\mathbf{0}$, thus $\mathbf{M}^{2}=-\det(\mathbf{M})\mathbf{I}=-\Omega^{2}{\bf I}$. Note that in oscillatory stable regime, $\det(\mathbf{M})=(-i(\Delta+2U|\Psi_{SS}|^{2}))(i(\Delta+2U|\Psi_{SS}|^{2}))-(-iU\Psi_{SS}^{2})(iU(\Psi_{SS}^{*})^{2})=(\Delta+2U|\Psi_{SS}|^{2})^{2}-|U\Psi_{SS}^{2}|^{2}=\Omega^{2}\geq 0$.
@@ VII Truncated Wigner Approximation (TWA)
For a symmetric vacuum state, the variance of the complex variable is $\langle|\alpha|^{2}\rangle=\langle X^{2}\rangle+\langle P^{2}\rangle=1/2$, with $\langle X^{2}\rangle=\langle P^{2}\rangle=1/4$ and $\Delta X=\Delta P=1/2$ (which is possibel only for the vacuum or a coherent/unsqueezed state). The product of uncertainties in linear system with Gaussian state is $\Delta X\Delta P=\frac{1}{4}$, corresponds to the minimal uncertainty state and minimal variance ($1/2$) in TWA. This is consistent with above symmetric ordering requirement $\overline{|\alpha|^{2}}=\langle\hat{n}\rangle+1/2$. The noise $\sqrt{\gamma/2}dW$ which inject randomness is isotropic while the nonlinearity ($U$) is anisotropic and cause non-Gaussian state with $\Delta X\Delta P>\frac{1}{4}$. As the system relaxes from the vacuum to a driven NESS, the total variance $|\alpha|^{2}$ increases due to the fluctuation part of TWA which suppressing the coherent amplitude and restoring the $U(1)$ symmetry for a local observer ($\langle\hat{a}_{i}\rangle=0$). While the sum of variances is bounded by the noise term, the individual quadrature variance (e.g., $\Delta X^{2}$) can be compressed below 1/4.
The Kerr term $\hat{n}(\hat{n}-1)$ maps to the Wigner symbol $H_{int}=\frac{U}{2}(|\alpha|^{4}-2|\alpha|^{2}+1/2)$. The drift $\mathcal{A}$ in Langevin equation is derived from the derivative (Poisson bracket) $\dot{\alpha}=-i\frac{\partial H_{int}}{\partial\alpha^{*}}$,
@@
where the quantum bias 1-1 is related to the above bias 1/2 in the operator $\alpha$ in symmetric (Weyl) ordering and ensures the classical trajectories obeys the quantum commutation relations during evolution. To find $(\hat{a}^{\dagger}\hat{a}\cdot\hat{a})_{W}$, we apply the Moyal product
$(|\alpha|^{2}-1/2)\star\alpha=(|\alpha|^{2}-1/2)\alpha+\frac{1}{2}\left(\frac{\partial(|\alpha|^{2}-1/2)}{\partial\alpha}\frac{\partial\alpha}{\partial\alpha^{*}}-\frac{\partial(|\alpha|^{2}-1/2)}{\partial\alpha^{*}}\frac{\partial\alpha}{\partial\alpha}\right)=(|\alpha|^{2}-1)\alpha$ (86)
where $\frac{\partial\alpha}{\partial\alpha^{*}}=0$, $\frac{\partial(|\alpha|^{2}-1/2)}{\partial\alpha^{*}}=\frac{\partial(\alpha\alpha^{*}-1/2)}{\partial\alpha^{*}}=\alpha$. The term $U(|\alpha|^{2}-1)\alpha$ in the Langevin equation is the exact Weyl symbol of the quantum operator $U\hat{a}^{\dagger}\hat{a}\hat{a}$. The 1-1 is a quantum correction arising from the non-commutativity of the bosonic operators, ensuring that the semiclassical trajectories correctly account for the underlying quantum statistics.
@@
Note that in any cases, we consider the deturning term $\propto\Delta n<0$ (i.e., $i\Delta\alpha$ in Heisenberg equation of motion), in consistent with the content in above sections. From the Heisenberg equation $\dot{\hat{a}}=-i[\hat{a},\hat{H}$], the interaction part is $i[\hat{a},\frac{U}{2}\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}]=-iU\hat{a}^{\dagger}\hat{a}\hat{a}=-iU\hat{n}\hat{a}$. The TWA variables $\alpha$ in symmetric (Weyl) ordering can be used to map the operator product $\hat{n}\hat{a}=\hat{a}^{\dagger}\hat{a}\hat{a}$ to phase space, by the Weyl symbol of $\hat{a}^{\dagger}\hat{a}\hat{a}$ is $|\alpha|^{2}\alpha-\alpha=(|\alpha|^{2}-1)\alpha$, where 1-1 is a quantum correction (ordering bias) that accounts for the commutation relation $\hat{a},\hat{a}^{\dagger}]=1$. The term $|\alpha|^{2}\alpha$ represents the classical mean-field interaction. The term $\alpha$ represents the self-interaction with vacuum noise that subtracts the over-counted zero-point energy inherent in the Wigner representation. The hopping term $J(\alpha_{l-1}+\alpha_{l+1})$ describes the coherent exchange of particles between neighboring sites in the chain. The dissipative damping $\frac{\gamma_{l}}{2}\alpha_{l}$ comes from the $\mathcal{D}[\sqrt{\gamma}\hat{a}$] term in the Lindblad equation. It causes the amplitude to decay exponentially toward zero in the absence of a drive. The stochastic term represents the vacuum noise entering through the loss channels, with the Markovian white noise $\langle d\mathcal{W}_{l}(t)d\mathcal{W}_{k}^{*}(t^{\prime})\rangle=\delta_{lk}\delta(t-t^{\prime})dt$. Since the loss is local to each site, $\mathcal{B}$ is a diagonal matrix with $\mathcal{B}_{ll}=\sqrt{\gamma_{l}/2}$ and scales the random noise by the square root of the dissipation rate. The complex wiener process $d\mathcal{W}_{l}$ represents independent random variables sampled at each time step. The evolution of the fluctuation spinor $\mathbf{v}=(\delta\hat{a},\delta\hat{a}^{\dagger})^{T}$ is governed by the linearized equation $\dot{\mathbf{v}}=\mathbf{L}_{\text{eff}}\mathbf{v}$. The matrix $\mathbf{L}_{\text{eff}}$ is defined as the Jacobian matrix of the total deterministic drift function $\dot{\alpha}=\mathcal{A}(\alpha,\alpha^{*})$,
@@
Figure 4: Distribution of final-time amplitudes $\alpha_{l}(t)$ across the trajectory ensemble in phase space. (a) The dephasing nonlinear regime ($U=2.5,F=6,J=0.8$) where nonlinearity introduces phase scrambling. This is a intermediate phase between the linear phase and chaotic regime, which can also be called as called symmetric regular phase or dissipatively stabilized phase. At site 1, the drive $F$ is strong enough to maintain a coherent offset. At sites 2 and 3, the scrambling is stronger than the coherence transfer $J$ and cause strong dephasing where the operator OTOC decays to zero at long time by dissipation and by phase averaging This generates diffusion and randomizes the phase, and the Wigner distribution is centered at (0,0), effectively restoring the $U(1)$ symmetry. The nonlinear phase-diffusive ring distribution can be observed only at site 2 and site 3. (b) The linear regime ($U=0,F=1,J=0.8$) where all sites are displaced from the origin. The phase is pinned and U(1) symmetry is broken by the drive because the coherence $J$ is dominate over the randomizing force. Site 2 and site 3 signify the thermal state with small local chemical potential such that the photons are not numerous enough to overcome vacuum fluctuations, so the distribution is centered at the origin and filled in by noise. This explain the absence of photon blockade and the scrambled prethermal phase.
Figure 5: Phase-space distributions ($\text{Re }\alpha$ vs. $\text{Im }\alpha$) for a three-site driven-dissipative Bose-Hubbard chain under noiseless evolution. (a) shows the response under a weak coherent drive ($F=1.0$), while (b) represents a strong drive ($F=6.0$). Blue points represent the ensemble endpoint distribution of 1,000 trajectories at $t_{\text{max}}=50$, each initialized with random vacuum Wigner fluctuations to simulate initial quantum uncertainty. Red trajectories illustrate the deterministic path of a single initial condition. At weak drive, site 1 exhibits a diffused, phase-nonuniform limit cycle where nonlinearity shears the initial Wigner phase uncertainty into a broad arc. Site 3 remains a stable fixed-point near the origin due to high local dissipation. At strong drive, site 1 converges to a stable fixed-point attractor. Site 2 displays a deterministic, phase-nonuniform limit cycle, appearing as a sharp blue crescent due to the competition between nonlinear phase-shearing and the coherent phase-locking of the drive. The lack of a closed red loop at site 2 suggests a slow approach to the limit-cycle manifold or a period longer than the sampling window.
In Fig.5, we set the dynamical noise as zero. The cloud-like distribution vanish in the absence of noise-driven phase diffusion, where there is only the initial noise from the vacuum fluctuations (i.e., deterministic evolution). The blue points show an steady-state ensemble snapshot distribution at long time, where we consider 1000 trajectories each corresponds to a unique initial condition. The spread of blue points reflects phase diffusion and noise-induced mixing in the stochastic semiclassical framework. It indicates loss of phase locking (phase becomes uncertain) and increased fluctuations. The red points show a single-run time-parametrized path in phase space (a time series from one single initial condition). This proves that the decay is due to the classical phase diffusion and damping instead of deterministic chaos. Thus our result shows that the random initial phases and dynamical noise cannot guarantees the uniform phase mixing and hence a uniformly filled ring by itself.
Site 2 for $F=6.0$ exhibits a phase-nonuniform attractor due to the phase-locking. The increased field amplitude $|\alpha|\propto F/\gamma$ strengthens the deterministic limit-cycle dynamics[26] where system’s amplitude and phase oscillate deterministically and phase along the cycle depends on the random initial condition. For $F=1$ at site 1, there is a diffused phase-nonuniform limit cycle. Here the spreading arises from nonlinearity shearing of the initial Wigner phase uncertainty by $U|\alpha|^{2}$ (deterministic mapping of initial quantum fluctuations into a nonlinear phase space) as well as the finite-time deterministic sampling[27], rather than environmental noise. For $F=6$ at site 1, the red trajectory spirals toward a stable fixed point attractor (equilibrium steady state). For $F=1.0$ at site 2, there is a diffused ring (partially phase-locked) where the phase is relatively undetermined. Because the drive is weaker, the phase-locking is less dominant, allowing the noise to spread the points more effectively around the circle. The spread of the noisy (TWA/Langevin) ensemble around the deterministic attractor quantifies noise-induced diffusion phase diffusion. In a semiclassical mapping, this diffusion is related to the loss of phase coherence, but it should not be claimed as a direct, quantitative decoherence measure without additional checks.
The nonlinearity $U$ causes different amplitudes to rotate at different angular velocities, stretching the initial Wigner distribution into a ring or arc. The larger dissipation at site 3 suppresses the amplitude and operator growth, results in a small phase-nonuniform limit-cycle (large $F$) or cluster (small $F$). Transition from $F=1$ to $F=6$ shows the system moving from a regime dominated by quantum fluctuations to one dominated by nonlinear dynamical instabilities ($U|\alpha|^{2}$) with phase nonuniform ring.
The drive ($F$) has a competing effect on the phase of the system. In one hand, it increases the steady-state amplitude $|\alpha|$. For nonlinear shift $U|\alpha|^{2}$, a larger $F$ indirectly increases the phase-shearing (different amplitudes rotating at different speeds) and spread the phase. In the other hand, $F$ is a coherent drive with a fixed phase and suppresses diffusion. This can be evidenced by considering the drive $F$ with randomized drive phase across trajectories. It randomize the phase to an extend independent of site. While larger nonlinearity $U$ will cause larger difference of $|\alpha|$ between sites.